### v1

# Improving the Cosmological Constraints by Inferring the Formation Channel of Extreme-mass-ratio Inspirals

Abstract
Extreme-mass ratio inspirals (EMRIs) could be detected by space-borne gravitational-wave (GW) detectors, such as the Laser Interferometer Space Antenna (LISA), TianQin and Taiji. In general, locating EMRIs by space-borne GW detectors can help us select the candidate host galaxies which can be used to infer the cosmic expansion history. In this paper, we show that the localization information of EMRIs can also be used to select the candidate host AGNs which can be used to infer the formation channel of the EMRIs and extract more precisely the redshift probability distributions of the EMRIs. Using the EMRIs that are expected to be detected by TianQin and LISA and the galaxy catalog that can be provided by the Chinese Space Station Telescope for the analysis, we find that TianQin can constrain the Hubble-Lemaître constant $H_{0}$ to a precision of about $3\%-8\%$ and the dark energy equation of state parameter $w_{0}$ to about $10\%-40\%$, the network composed of TianQin and LISA can improve the precisions of $H_{0}$ and $w_{0}$ to about $0.4\%-7\%$ and $4\%-20\%$, respectively, without considering the effects of AGNs. Furthermore, combining the detected EMRIs with AGN catalog by the statistical framework of likelihood-ratio-based, we find TianQin can establish the EMRI-AGN correlation with about $500$ EMRIs if all of the EMRIs indeed originate in AGNs, and the TianQin+LISA network can reduce this required number to about $30$. Once the EMRI-AGN correlation is confirmed, the combination of the EMRIs and AGN catalog can significantly improve the constraints on the cosmological parameters. These results demonstrate the potentials of using EMRIs as well as galaxy and AGN catalogs to constrain the cosmological parameters and astrophysical models.
Gravitational waves (678), Black holes (162), Sky surveys (1464), Hubble constant (758), Bayesian statistics (1900), Poisson distribution (1898)
software: numpy (van der Walt et al., 2011), scipy (Virtanen et al., 2020), emcee (Foreman-Mackey et al., 2013, 2019), matplotlib (Hunter, 2007), corner (Foreman-Mackey, 2016), GetDist (Lewis, 2019), and seaborn (Waskom, 2021). thanks: Boya fellow
1 Introduction
The first three observing runs of the Advanced LIGO and Virgo, and KAGRA detectors (LVK) have detected about $90$ gravitational-wave (GW) events triggered by the mergers of compact-object binaries (Abbott et al., 2016, 2017a, 2019, 2021b, 2023a). In addition, increasing evidence for the existence of GW signal in the nanohertz band is accumulating thanks to the observations of pulsar timing arrays (Agazie et al., 2023; Antoniadis et al., 2023; Reardon et al., 2023; Xu et al., 2023), although the origin of this GW signal has yet to be identified. These observations opened a new era of observational astronomy, in which GWs can be used as new probes of fundamental physics (Abbott et al., 2021c, d), astrophysics (Fishbach & Holz, 2017; Abbott et al., 2020b; Fishbach et al., 2021; Abbott et al., 2023b; Bartos et al., 2017; Veronesi et al., 2022, 2023), and the cosmic expansion history (Abbott et al., 2017b; Soares-Santos et al., 2019; Fishbach et al., 2019; Hotokezaka et al., 2019; Palmese et al., 2020; Vasylyev & Filippenko, 2020; Abbott et al., 2021a; Wang & Giannios, 2021; Finke et al., 2021; Abbott et al., 2023c).
The idea of combining a GW event and its electromagnetic (EM) counterpart to establish a “standard siren” which can be used to probe the expansion history of the Universe, such as measuring the Hubble-Lemaître constant, was first proposed by Schutz (1986). Many researchers followed this idea and have established the methodological framework even before the detection of GW (Marković, 1993; Holz & Hughes, 2005; Dalal et al., 2006; Arun et al., 2009; Petiteau et al., 2011; Babak et al., 2011; Taylor et al., 2012; Del Pozzo, 2012; Nissanke et al., 2013). The first successful measurement of the Hubble-Lemaître constant using standard siren was accomplished in 2017 (Abbott et al., 2017b) immediately after the detection of a binary neutron star merger in both GWs (Abbott et al., 2017a) and EM waves (Abbott et al., 2017c). Even when a unique EM counterpart is difficult to identify, the redshift may be inferred from the statistical distribution of the galaxies in the same sky area localized by GW detectors (Schutz, 1986; MacLeod & Hogan, 2008; Petiteau et al., 2011; Del Pozzo, 2012). Such “dark sirens” also provide important constraints on the Hubble-Lemaître constant, e.g., in the cases of GW170814 (Soares-Santos et al., 2019) and GW190814 (Palmese et al., 2020; Vasylyev & Filippenko, 2020).
So far, using the third Gravitational-wave Transient Catalog (GWTC-3), the LVK collaboration have measured the Hubble-Lemaître constant to a precision of $10\%$ (Abbott et al., 2023c, without using the information from the afterglow of GW170817). Better constraint can be achieved by the next-generation ground-based GW detectors, such as the Advanced LIGO A+/Voyager (Abbott et al., 2020a), Virgo Plus (Acernese et al., 2023), the Einstein Telescope (ET) (Punturo et al., 2010), and Cosmic Explorer (CE) (Reitze et al., 2019). Then we can infer not only the cosmic expansion history, but also the fractional density parameters of the Universe and the equation of state (EoS) of the dark energy (Del Pozzo et al., 2017; Cai & Yang, 2017; Zhao & Wen, 2018; Belgacem et al., 2019; Jin et al., 2020; Yu et al., 2020; You et al., 2021; Chen et al., 2021; Bonilla et al., 2022; Leandro et al., 2022; Song et al., 2022). These measurements will help solve the most difficult questions in today’s cosmology, such as the inconsistency between the measured values of the Hubble-Lemaître constant in the early and late universe (Feeney et al., 2019; Borhanian et al., 2020; Califano et al., 2023; Gupta, 2023), which is known as the “Hubble tension” (Freedman, 2017; Aghanim et al., 2020; Riess et al., 2021; Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022).
In about a decade, we are likely to have space-borne GW detectors, such as the Laser Interferometer Space Antenna (LISA) (Amaro-Seoane et al., 2017; Colpi et al., 2024), TianQin (Luo et al., 2016; Mei et al., 2021), or Taiji (Hu & Wu, 2017). These detectors are sensitive to the GWs in the band of $0.1~{}\!{\rm mHz}-1~{}\!{\rm Hz}$. In this band, the potential GW sources which can be used as standard sirens include stellar-mass ($\sim 10M_{\odot}$) binary black holes (BHs), massive BH (MBH) binaries ($10^{5}-10^{8}M_{\odot}$), and extreme-mass-ratio inspirals (EMRIs) (see Auclair et al., 2023, for a review). The usage of the first two sources in cosmology have been extensively studied (Holz & Hughes, 2005; Petiteau et al., 2011; Tamanini et al., 2016; Caprini & Tamanini, 2016; Del Pozzo et al., 2018; Wang et al., 2022, 2022a; Zhu et al., 2022a, b; Muttoni et al., 2022; Jin et al., 2023). These studies show that the stellar-mass binary BHs detectable by space-borne GW detectors are predominately in the local universe (Sesana, 2016; Kyutoku & Seto, 2016; Liu et al., 2020), while MBH binaries (MBHBs) are among the loudest GW sources in the milli-Hertz band and hence can be detected to high redshift ($z\gtrsim 10$, Holz & Hughes, 2005; Cutler & Vallisneri, 2007; Klein et al., 2016; Amaro-Seoane et al., 2017; Wang et al., 2019a).
The third type of low-frequency GW sources which can be used as standard sirens are EMRIs. EMRI normally resides at the center of a galaxy and consists of an MBH and an inspiralling stellar-mass BH (Amaro-Seoane, 2018). The potential of using LISA EMRIs to study cosmology was first pointed out in MacLeod & Hogan (2008), and later studied in more detail by the literatures (Laghi et al., 2021; Liu et al., 2023). These earlier studies considered only the EMRIs with $z<1$ where the galaxy catalogs are relatively more complete. However, population studies have shown that space-borne GW detectors can catch EMRIs up to $z\approx 3$ or higher (Babak et al., 2017; Fan et al., 2020; Bonetti & Sesana, 2020). At such high redshifts, although the current galaxy catalog is incomplete, future space telescopes will be able to conduct a more complete survey of the galaxies (Gong et al., 2019; Ivezić et al., 2019; Scaramella et al., 2022). In addition, beside LISA, other space-borne GW detectors, such as TianQin and Taiji, can also detect EMRIs. A joint observation by multiple detectors in principle can improve the sky localization of a GW source by orders of magnitude (Ruan et al., 2020; Gong et al., 2021). These observational prospects in general should improve the constraint on cosmological parameters by EMRIs.
The prospect of using EMRIs to constrain the cosmological parameters relies on not only the completeness of a galaxy catalog, but also the accuracy of the redshift which can be inferred from the galaxy catalog. When extracting the redshift probability distributions of EMRIs from galaxy catalogs, previous works (MacLeod & Hogan, 2008; Laghi et al., 2021; Liu et al., 2023) have assumed that all galaxies have an equal probability to host an EMRI. However, in theory EMRIs could form in either normal galaxies (Amaro-Seoane, 2018) or active galactic nuclei (AGNs, e.g. Yunes et al., 2011), and recent models suggest that the latter channel could be more important because the accretion disks in AGNs can capture and drive a large number of stellar-mass BHs to migrate towards the center (Pan et al., 2021). Therefore, what type of galaxies in a catalog can be considered as the hosts of EMRIs remains a question.
Recently, a statistical framework has been proposed which allows one to test the spatial correlation between GW sources and AGNs even in the absence of identifying the host galaxies of the GW sources Bartos et al. (2017). The framework utilizes a Poisson distribution to describe the likelihood of the number of candidate host AGNs for the GW source, and constrains the spatial correlation between the GW sources and AGNs by a likelihood-ratio-based method. Several works have demonstrated the effectiveness of this framework in identifying the host galaxies of the LVK binary BHs (Veronesi et al., 2022, 2023) or the MBHBs of future space-borne GW detectors (Zhu & Chen, 2024). Such a method in principle can be used to infer the fraction of EMRIs which originate in AGNs. Knowing this fraction will help us better weigh the large number of candidate host galaxies in the sky area localized for an EMRI, so that we can more precisely extract the redshift probability distributions and measure cosmological parameters.
Therefore, in this work we will apply the aforementioned statistical framework to EMRIs, and study its effectiveness in constraining the cosmological parameters and testing the spatial correlation between EMRIs and AGNs. To facilitate the study, we consider TianQin, as well as a network composed of TianQin and LISA, as the future observational instruments. This paper is organized as follows. In Section 2 we introduce the statistical analysis framework that we use to infer the cosmic expansion history and test the EMRI-AGN correlation. In Section 3 we conduct mock observations. The results are shown in Section 4 for the cosmological prospects and Section 5 for the prospects of inferring the EMRI-AGN correlation. In Section 6, we discuss some of the factors that can affect the resuls. Finally, in Section 7 we summarize the main findings of this work and the implications of our results.
2 Methodology
2.1 Cosmological model
We assume that the Universe is flat (Aghanim et al., 2020; Dhawan et al., 2021) and can be described by the Friedmann-Lemaître-Robertson-Walker metric. From the Friedmann equations, the expansion rate of the Universe can be described by
$\!\!\!\!\!\!\!\!\!\!\!\!\!H(z)=H_{0}\sqrt{\Omega_{M}(1\!+\!z)^{3}+\Omega_{\Lambda}\exp\!\left[3\!\!\int_{0}^{z}\!\frac{1\!+\!w(z^{\prime})}{1\!+\!z^{\prime}}\mathrm{d}z^{\prime}\right]},$ (1)
where $H(z)$ is called as the Hubble-Lemaître parameter and $H_{0}\equiv H(z=0)$ is its current value, $\Omega_{M}$ and $\Omega_{\Lambda}$ are respectively the ratios of the total matter density and dark energy density to the critical density of the Universe (they satisfy $\Omega_{M}+\Omega_{\Lambda}=1$ in a flat-universe), and $w(z)$ describes the dark energy EoS as a function of redshift $z$.
In this work, we consider two dark energy models, namely the standard model—the cosmological constant $\Lambda$, and the Chevallier-Polarski-Linder (CPL) model (Chevallier & Polarski, 2001; Linder, 2003). (i) For the cosmological constant model, the EoS of the dark energy is constantly equal to minus one, i.e., $w(z)\equiv-1$. This model plus the cold dark matter model, plus the spatially-flat assumption, make up the so-called standard cosmological model—flat-$\Lambda$CDM. (ii) For the CPL model, the EoS of the dark energy is modeled as a function, $w(z)=w_{0}+w_{a}{z}/{(1+z)}$ (Chevallier & Polarski, 2001), here $w_{0}$ and $w_{a}$ are the extra parameters. The analytic expression of Equation (1) is $H(z)=H_{0}\sqrt{\Omega_{M}(1\!+\!z)^{3}+\Omega_{\Lambda}(1\!+\!z)^{3(1+w_{0}+w_{a})}\exp\!\big{(}\!\!-\!\frac{3w_{a}z}{1+z}\!\big{)}}$. In particular, when $w_{0}=-1$ and $w_{a}=0$, the CPL model will degenerate with the cosmological constant model.
Regardless of dark energy models, we have the “luminosity distance--redshift” ($D_{L}-z$) relation under a flat-universe according to the expansion effect of the Universe, i.e.,
$D_{L}=c(1+z)\!\int_{0}^{z}\!\frac{1}{H(z^{\prime})}\mathrm{d}z^{\prime},$ (2)
where $c$ is the speed of light. In our simulations, we fix all the cosmological parameters to the values: $H_{0}\equiv h\times 100~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}=67.8~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$ (here $h$ is the dimensionless Hubble-Lemaître constant), $\Omega_{M}=0.307$, $w_{0}=-1$, and $w_{a}=0$.
2.2 Bayesian framework for GW cosmology
To constrain the cosmological parameters by combining the GW and EM data, we adopt the Bayesian framework presented in Chen et al. (2018). However the estimations of the completeness of the galaxy catalogs were modified according to the correlation between the EMRIs’ MBHs and their host galaxies, such as the approximate linear relationship between the masses of MBHs and the (bulge) luminosities of the galaxies (which is known as the $M_{\rm MBH}-L$ relation; Ferrarese & Merritt, 2000; Scott & Graham, 2013; Kormendy & Ho, 2013).
Given a GW data set composed of $N$ independent GW events $\mathcal{D}^{\rm gw}\equiv\{d^{\rm gw}_{1},d^{\rm gw}_{2},\dots,d^{\rm gw}_{i},\dots,d^{\rm gw}_{N}\}$ as well as the corresponding data set of EM counterparts $\mathcal{D}^{\rm em}\equiv\{d^{\rm em}_{1},d^{\rm em}_{2},\dots,d^{\rm em}_{i},\dots,d^{\rm em}_{N}\}$, the posterior probability distribution of the cosmological parameters (denote as ${\bf\Omega}\equiv\{H_{0},\Omega_{M},w_{0},w_{a}\}$) that we are interested in can be expressed as
$\!\!\!\!\!\!\!p({\bf\Omega}|\mathcal{D}^{\rm gw},\mathcal{D}^{\rm em},I)\propto p_{0}({\bf\Omega}|I)p(\mathcal{D}^{\rm gw},\mathcal{D}^{\rm em}|{\bf\Omega},I)$
$\propto p_{0}({\bf\Omega}|I)\prod_{i}p(d^{\rm gw}_{i},d^{\rm em}_{i}|{\bf\Omega},I)$ , (3)
where $I$ indicates all the relevant additional information, $p_{0}({\bf\Omega}|I)$ is the prior probability distribution for ${\bf\Omega}$, and $p(d^{\rm gw}_{i},d^{\rm em}_{i}|{\bf\Omega},I)$ represents the likelihood of observing the $i$th GW event as well as the corresponding EM signal. Because the data contain noise such that different sources with different right ascensions $\alpha$ and declinations $\delta$, luminosity distances $D_{L}$, redshifts $z$, masses $M$, and luminosities $L$ may generate the same observational data. Therefore, the likelihood needs to marginalize all possible $(\alpha,\delta,D_{L},z,M,L)$, i.e.,
$p(d^{\rm gw}_{i},d^{\rm em}_{i}|{\bf\Omega},I)$
$\frac{\int\!p(d^{\rm gw}_{i}\!,\!d^{\rm em}_{i}\!,\!\alpha,\!\delta,\!D_{L},\!z,\!M,\!L|{\bf\Omega},\!I)\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}D_{L}\mathrm{d}z\mathrm{d}M\mathrm{d}L}{\beta({\bf\Omega}|I)},$ (4)
where $\beta({\bf\Omega}|I)$ is a normalization factor that use for accounts for the selection effects in GW and EM observations (Mandel et al., 2019; Chen et al., 2018; Soares-Santos et al., 2019).
To calculate the numerator in Equation (2.2), we separate $p(d^{\rm gw}_{i}\!,d^{\rm em}_{i}\!,\alpha,\delta,D_{L},z,M,L|{\bf\Omega},I)$ into
$p(d^{\rm gw}_{i},d^{\rm em}_{i},\alpha,\delta,D_{L},z,M,L|{\bf\Omega},I)$
$~{}p(d^{\rm gw}_{i},d^{\rm em}_{i}|\alpha,\delta,D_{L},z,M,L,{\bf\Omega},I)$
×$\!\!\times p_{0}(\alpha,\delta,D_{L},z,M,L|{\bf\Omega},I)$
$~{}p(d^{\rm gw}_{i}|\alpha,\delta,D_{L},M,{\bf\Omega},I)p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$
×$\!\!\times p_{0}(D_{L}|z,{\bf\Omega},I)p_{0}(M|z,L,{\bf\Omega},I)p_{0}(\alpha,\delta,z,L|{\bf\Omega},I).$ (5)
The term $p(d^{\rm gw}_{i}|\alpha,\delta,D_{L},M,{\bf\Omega},I)$ is the likelihood of the GW data (Finn, 1992; Sathyaprakash & Schutz, 2009) after marginalized over all other parameters that unrelated to the cosmic expansion history, and $p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$ is the likelihood of the corresponding EM data. For dark sirens, because there is no information from the observations of the EM counterparts, we can define $d^{\rm em}_{i}$ to null, as well as define the likelihood $p(d^{\rm em}_{i}|\alpha,\delta,z,L,{\bf\Omega},I)$ to constant (Chen et al., 2018). The term $p_{0}(D_{L}|z,{\bf\Omega},I)$ is a Dirac delta function, i.e., $p_{0}(D_{L}|z,{\bf\Omega},I)=\delta_{\rm D}\big{(}D_{L}-\hat{D}_{L}(z,{\bf\Omega})\big{)}$ (here $\hat{D}_{L}(z,{\bf\Omega})$ represents Equation (2)), because $D_{L}$ is a function of $z$ given a cosmology.
The term $p_{0}(M|z,L,{\bf\Omega},I)$ can be approximated as a uniform distribution of $M$, i.e.,
$p_{0}(M|z,L,{\bf\Omega},I)\propto$
$\mathcal{U}\big{[}(1\!+\!z)(\hat{M}(L)\!-\!3\sigma_{\!M}),(1\!+\!z)(\hat{M}(L)\!+\!2\sigma_{\!M})\big{]}(M),$ (6)
where the two values in square bracket $\mathcal{U}[...~{}\!,...]$ represent the lower and upper boundaries of the uniform distribution, $\hat{M}(L)$ represents the mass of the MBH as a function of the luminosity of the host galaxy, and $\sigma_{\!M}$ is the standard deviation. The expression of $p_{0}(M|z,L,{\bf\Omega},I)$ holds just for EMRI (and MBHB) GW sources due to the $M_{\rm MBH}-L$ relation, the reason for adopting a uniform rather than a Gaussian distribution is that the data for which the $M_{\rm MBH}-L$ relation is derived have many outliers (Ferrarese & Merritt, 2000; Scott & Graham, 2013; Kormendy & Ho, 2013). The function $\hat{M}(L)$ can be written as
$\!\!\!\!\!\!\!\!\!\!\!\!\!\lg\!\left(\!\frac{\hat{M}(L)}{M_{\odot}}\!\right)\!\approx a_{\rm fit}+b_{\rm fit}\!\left[\!M_{\odot}^{*}-2.5\lg\!\left(\!\frac{L}{L_{\odot}}\!\right)+23.4\right]\!,$ (7)
(Scott & Graham, 2013), where $a_{\rm fit}$ and $b_{\rm fit}$ are the fitting zeropoint and slope, and $M_{\odot}\!=\!1.989\times 10^{30}~{}\!{\rm kg}$, $M^{*}_{\odot}\!=\!+4.8\!~{}{\rm mag}$ and $L_{\odot}\!=\!3.8\times 10^{33}\!~{}{\rm erg~{}s}^{-1}$ are the mass, absolute magnitude, and luminosity of the Sun, respectively. The best-fitting parameters for Equation (7) in the $K$-band are $a_{\rm fit}=8.04$ and $b_{\rm fit}=-0.48$ with a root-mean-square scatter of $\sigma_{M}=0.4$ in $\lg(M/M_{\odot})$ direction (Scott & Graham, 2013). The reason, moreover, that the uniform distribution we adopt is not symmetric around $\hat{M}(L)$ is that, there is a fraction of galaxies with lighter MBHs compared to the predicted $\hat{M}(L)$ (Jiang et al., 2011; Scott & Graham, 2013; Kormendy & Ho, 2013).
The last term $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$ in Equation (2.2) is a prior distribution on the spatial positions and luminosities of galaxies. To get a proper prior, we adopt the standard assumption that the spatial probability distribution of compact binary GW sources is proportional to the number density of galaxies (Schutz, 1986; MacLeod & Hogan, 2008; Petiteau et al., 2011; Del Pozzo, 2012; Chen et al., 2018; Fishbach et al., 2019; Soares-Santos et al., 2019; Gray et al., 2020; Finke et al., 2021; Abbott et al., 2023c), while taking into account the selection effects that many galaxies are too dim to be recorded by surveys, we can express the prior as
$\!\!\!\!\!\!\!p_{0}(\alpha,\delta,z,L|$ ${\bf\Omega},I)=~{}f_{\rm compl}~{}\!p_{\rm obs}(\alpha,\delta,z,L|{\bf\Omega},I)$
$+(1-f_{\rm compl})~{}\!p_{\rm miss}(\alpha,\delta,z,L|{\bf\Omega},I),$ (8)
where $p_{\rm obs}(\alpha,\delta,z,L|{\bf\Omega},I)$ represents a distribution function of observed galaxies from EM surveys, $f_{\rm compl}$ is the completeness fraction of the observed galaxy catalog, and $p_{\rm miss}(\alpha,\delta,z,L|{\bf\Omega},I)$ represents the probability distribution function of the unobservable galaxies.
To calculate $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$, we need specific expressions for the two terms on the right-hand side of Equation (2.2). (i) The distribution function of observed galaxies $p_{\rm obs}(\alpha,\delta,z,L|{\bf\Omega},I)$ can generally be written as
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!p_{\rm obs}(\alpha,\delta,z,L|{\bf\Omega},$ $I)\propto\sum_{j=1}^{N_{\rm gal}}\Big{[}\delta_{\rm D}(\alpha-\alpha_{j})~{}\!\delta_{\rm D}(\delta-\delta_{j})$
×$~{}~{}\times\!\mathcal{N}[\bar{z}_{j},\sigma_{z;j}](z)~{}\!\mathcal{N}[\bar{L}_{j},\sigma_{L;j}](L)\Big{]},$ (9)
where $N_{\rm gal}$ is the total number of the observed galaxies and two $\mathcal{N}[\bar{x},\sigma_{x}](x)$ represent Gaussian distributions of $x$ with mean values $\bar{x}$ and standard deviations $\sigma_{x}$. Here, because the errors of sky positions of galaxies in EM surveys are much smaller than the sky localization errors of the GW sources, the distribution of $(\alpha,\delta)$ of galaxies can be approximately expressed as two Dirac delta functions. (ii) For the unobservable galaxies, their probability distribution function can be expressed as
$\!\!\!\!\!\!\!\!\!\!\!p_{\rm miss}(\alpha,\delta,z,L|{\bf\Omega},I)\propto{\rm erfc}(L)~{}\!\Phi(L,z)\frac{\mathrm{d}V_{\rm c}}{\mathrm{d}z\mathrm{d}\hat{\Omega}}\mathrm{d}z\mathrm{d}\hat{\Omega},$ (10)
where
$\!\!\!\!\!\!\!\!\!\!{\rm erfc}(L)=\frac{1}{\sqrt{2\pi\bar{\sigma}_{L}^{2}}}\!\int_{L}^{\infty}\!\!\exp\!\Bigg{[}\!\!-\frac{\big{(}L^{\prime}-\hat{L}_{\rm thr}(z,{\bf\Omega})\big{)}^{2}}{2~{}\!\bar{\sigma}_{L}^{2}}\Bigg{]}\!\mathrm{d}L^{\prime}$
is the complementary error function with mean value $\hat{L}_{\rm thr}(z,{\bf\Omega})$ and standard distribution $\bar{\sigma}_{L}$, $\Phi(L,z)$ represents the luminosity function of the galaxies at $z$ (Schechter, 1976; Dahlen et al., 2005; Marchesini et al., 2007), $V_{\rm c}$ is the comoving volume, and $\mathrm{d}\hat{\Omega}=\cos\delta\!~{}\mathrm{d}\alpha\mathrm{d}\delta$ represents the differential solid angle. The mean value $\hat{L}_{\rm thr}(z,{\bf\Omega})$ is a luminosity threshold for an observable galaxy with $z$ in a cosmology ${\bf\Omega}$, it depends on the limit magnitude $m_{\rm limit}^{*}$ of the galaxy survey, we can calculate it with
$\!\!\!\!\!\!\!\!\!\!\lg\!\left(\!\frac{\hat{L}_{\rm thr}(z,{\bf\Omega})}{L_{\odot}}\!\right)\!=$ $*\displaystyle~{}\!\frac{1}{2.5}\Bigg{[}M^{*}_{\odot}-m_{\rm limit}^{*}$
$+5\lg\!\left(\!\frac{\hat{D}_{L}(z,{\bf\Omega})}{1~{}\!{\rm Mpc}}\!\right)\!+25\Bigg{]}.$ (11)
The complete expression of $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$ additionally requires a reasonable estimation for the completeness of the observed galaxy catalog. In this work, we consider the role of the $M_{\rm MBH}-L$ relation for estimating the completeness of galaxy catalogs. We assume that the completeness depends not only on the galaxy catalog itself, but also on the property of the EMRI source (such as the mass of the MBH). The completeness fraction can be estimated by
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!f_{\rm compl}$ $\approx 1\!-\!\bigg{[}\!\int\!\!p(d^{\rm gw}_{i}|\alpha,\delta,D_{L},M,{\bf\Omega},I)p_{0}(D_{L}|z,{\bf\Omega},I)$
×$\times p_{0}(M|z,L,{\bf\Omega},I)$
×$\times p_{\rm miss}(\alpha,\delta,z,L|{\bf\Omega},I)\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}D_{L}\mathrm{d}M\mathrm{d}z\mathrm{d}L\bigg{]}.$ (12)
The $M_{\rm MBH}-L$ relation is the statistical basis that enables the term $p_{0}(M|z,L,{\bf\Omega},I)$ to connect the EMRI GW source to the galaxy catalog. From Equation (2.2), we can find that for an EMRI source with a more massive MBH, the completeness of the same galaxy catalog is higher.
Finally, to calculate the denominator $\beta({\bf\Omega}|I)$ in Equation (2.2), we adopt the same estimated approach as presented in Chen et al. (2018) (also see Gray et al., 2020; Finke et al., 2021; Abbott et al., 2023c; Zhu & Chen, 2023). For dark sirens, the denominator is determined by the GW data $d^{\rm gw}$ and the prior $p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)$, we have
$\!\!\!\!\!\!\!\!\!\!\!\!\beta({\bf\Omega}|I)$ $=\!\!\!\int\limits_{{\rm detectable}}\!\!\!\!\!\!p(d^{\rm gw}|\hat{D}_{L}(z,{\bf\Omega}),M,I)p_{0}(M|z,L,{\bf\Omega},I)$
×$\!\!\!\times p_{0}(\alpha,\delta,z,L|{\bf\Omega},I)~{}\!\mathrm{d}d^{\rm gw}\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}M\mathrm{d}z.$ (13)
Here, the equation represents to integrate over the possible GW data $d^{\rm gw}$ that the signal-to-noise ratio (SNR) is greater than the detection threshold.
2.3 Statistical framework for constraining the formation channel of EMRIs
An effective constraint on the formation channel of EMRIs can be obtained by testing the spatial correlation between EMRIs and AGNs. We set the null hypothesis of our test to be that EMRIs can reside in either normal galaxies or AGNs. It means that the probability that a detected EMRI originates in an AGN is proportional to the fraction of AGNs among all galaxies. Our alternative hypothesis is that EMRIs predominantly reside in AGNs.
To statistically test our null hypothesis, we use the same statistical framework presented in Zhu & Chen (2024), which was developed upon the work of Bartos et al. (2017) and modified to test the spatial correlation between MBHB mergers and AGNs. Given that the spatial number density of AGNs is $\rho_{\rm agn}$, and the spatial localization error volume of the $i$th detected EMRI is $\Delta V_{i}$, in the null hypothesis the number $N_{{\rm agn},i}$ of AGNs within the error volume $\Delta V_{i}$ will follow a Poisson distribution with an expectation of $\rho_{\rm agn}\Delta V_{i}$, i.e.,
$B_{i}(N_{{\rm agn},i})={\rm Poisson}\Big{(}N_{{\rm agn},i},\rho_{\rm agn}\Delta V_{i}\Big{)}.$ (14)
In the alternative hypothesis, there is one guaranteed host AGN within $\Delta V_{i}$. Therefore, the number of the rest non-host AGNs, $N_{{\rm agn},i}-1$, will follow the Poisson distribution with the expectation $\rho_{\rm agn}\Delta V_{i}$, i.e.,
$S_{i}(N_{{\rm agn},i})={\rm Poisson}\Big{(}N_{{\rm agn},i}-1,\rho_{\rm agn}\Delta V_{i}\Big{)}.$ (15)
Taking into account our lack of knowledge about the formation channel of EMRIs, we assume that a fraction $f_{\rm agn}$ of detected EMRIs are from AGNs, while the rest detected EMRIs originate in normal galaxies. Effectively, it means that the probability that an individual EMRI originates in an AGN is equal to $f_{\rm agn}$. Furthermore, taking into account the incompleteness of the AGN catalog, we can write the individual likelihood of the $i$th detected EMRI as
$\!\!\!\!\!\!\!\!\mathcal{L}_{i}(f_{\rm agn})=f_{\rm agn}f_{{\rm compl},i}^{\rm agn}S_{i}+\Big{(}\!1-f_{\rm agn}f_{{\rm compl},i}^{\rm agn}\!\Big{)}B_{i}$ (16)
(Zhu & Chen, 2024), where $f_{{\rm compl},i}^{\rm agn}$ represents the completeness fraction of the AGN catalog relative to the $i$th detected EMRI. For the estimation of $f_{{\rm compl},i}^{\rm agn}$, we follow exactly the approach of Zhu & Chen (2024), by considering the correlation between the luminosities of AGNs and the masses of their central MBHs. For a GW catalog composed of $N$ EMRIs, the total likelihood can be expressed as
$\mathcal{L}(f_{\rm agn})=\prod_{i}\mathcal{L}_{i}(f_{\rm agn}).$ (17)
Finally, to statistically constrain the fraction $f_{\rm agn}$ in Equation (17), we follow Bartos et al. (2017) to introduce a likelihood ratio
$\lambda=2\ln\!\left[\frac{\mathcal{L}(f_{\rm agn})}{\mathcal{L}(0)}\right].$ (18)
In the null hypothesis, in which the value of $f_{\rm agn}$ equals to the fraction of AGNs among all galaxies in the Universe, the likelihood ratio $\lambda$ is governed by the random fluctuation of the spatial number density of AGNs, and will follow a “background distribution” $P_{\rm bg}(\lambda)$. A very important property is that the median of $\lambda$ derived from the detected EMRIs will deviate more from $P_{\rm bg}(\lambda)$ when the fraction of the EMRIs originate in AGNs is larger. In the analysis, $P_{\rm bg}(\lambda)$ can be determined by Monte-Carlo simulations, and the deviation of median $\lambda$ relative to $P_{\rm bg}(\lambda)$ can be quantified by an $p$-value. If this $p$-value becomes less than $0.00135$, we can claim that we can reject the null hypothesis with a significance of $3\sigma$, i.e., confirm the spatial correlation between the EMRIs and AGNs.
3 Mock Data
3.1 Population models of EMRIs
The EMRI rate is the primary feature of the EMRI population, which is mainly determined by four factors (Babak et al., 2017): (i) the properties of the MBH population, including the mass function and spin distribution of MBHs; (ii) the $M-\sigma$ relation, including the properties of the stellar didtribution surrounding the MBH, the presence and regrowth time of cusp erosion following MBHB formation and evolution, and the EMRI duty cycle; (iii) the ratio of plunges to EMRIs; (iv) the characteristic masses of inspiralling stellar-mass compact objects. All of these factors are uncertain in astrophysics, and different conditions produce different EMRI population models. In this work, we adopt 11 EMRI population models proposed in Babak et al. (2017), i.e., from M1 to M12 (except the M11 model, which predicts a very low EMRI rate so that LISA and TianQin detecting almost no EMRIs, e.g., Babak et al., 2017; Fan et al., 2020). We set the eccentricities of EMRIs at plunge to be uniformly distributed in the range from 00 to $0.2$. The main differences between the various models are summarized in the Table I of Babak et al. (2017).
3.2 Mock observations
To estimate the abilities of detectors in constraining the cosmic expansion history and the formation channel of EMRIs, we need to simulate (i) the detected EMRI catalogs, and (ii) the redshift probability distribution of the candidate host galaxies as well as (iii) the number of candidate host AGNs corresponding to each detected EMRI.
To generate the detected EMRI catalogs, we follow the approach of Babak et al. (2017) and Fan et al. (2020). First, we adopt a galaxy catalog from the MultiDark cosmological simulations (Klypin et al., 2016; Croton et al., 2016) to select the host galaxies of the EMRIs, with the mass of the MBH at the host galaxy center concordant with the mass of the MBH of the corresponding EMRI. Second, we generate the GW signals of EMRIs using the analytic kludge Kerr (AKK) waveform model (Barack & Cutler, 2004; Babak et al., 2017). The reason we adopt the AKK waveform model here is to facilitate comparison with the previous works (MacLeod & Hogan, 2008; Laghi et al., 2021), although several new waveform models (Chua et al., 2017, 2021; Hughes et al., 2021; Isoyama et al., 2022) have been proposed recently. Third, we calculate the detector responses to the EMRI GW signals, and consider a total of three detector configurations, namely
  • TianQin (TQ): the default situation in which three satellites form a constellation and operate in a “3 months on + 3 months off” mode, with a mission lifetime of five years (Luo et al., 2016);
  • TianQin I+II (TQ I+II): twin constellations of satellites with perpendicular orbital planes, that operate in a relay mode and can avoid the 3 months gap in data (Fan et al., 2020; Liang et al., 2022);
  • TianQin+LISA (TQ+LISA): a network composed of TianQin and LISA, the LISA configuration according to Amaro-Seoane et al. (2017). In this work, we assume that LISA and TianQin have a five year mission overlap and only consider EMRIs that are individually detectable by both TianQin and LISA.
Finally, we set the threshold of SNR for GW detection to $20$, and estimate the spatial localization errors of EMRIs with the Fisher information matrix (Cutler & Flanagan, 1994; Vallisneri, 2008).
Figure 1: Panels (a), (b), (c), and (d) show, respectively, the distributions of the total $D_{L}$ estimation errors, the sky localization errors, the numbers of the candidate host galaxies, and the expected numbers of the candidate host AGNs, for detectable EMRIs under the M1 population model. The colors indicate TianQin (black), TianQin I+II (blue), and TianQin+LISA (red). And in Panels (c) and (d), the solid and dashed lines, respectively, represent the conditions in which the cosmological parameters are fixed and freely varying within a prior range, when transforming the ranges of luminosity distances of EMRIs into the ranges of redshifts.
Here, we take the M1 model as the fiducial model as in Laghi et al. (2021). For example, the distribution of the spatial localization errors (including the $D_{L}$ estimation and sky localization errors) of various detector configurations are shown in Panels (a) and (b) of Figure 1. We find that the medians of the estimation errors for luminosity distance are about $7\%$, $6\%$, and $3\%$ for TianQin, TianQin I+II, and TianQin+LISA, respectively, after accounting for additional systematic errors due to the weak lensing effects (Hirata et al., 2010; Cusin & Tamanini, 2021) and peculiar velocities of the host galaxies (Kocsis et al., 2006). The corresponding medians errors for sky localization are about $2.6~{}\!{\rm deg}^{2}$, $0.8~{}\!{\rm deg}^{2}$, and $0.02~{}\!{\rm deg}^{2}$, respectively.
For the cosmological inferences, in order to simulate the redshift probability distributions of the detected EMRIs, we select their candidate host galaxies from the MultiDark galaxy catalog for which the selection effects of surveys have been considered. Here we set the limiting magnitude of the MultiDark galaxy catalog to $m_{\rm limit}^{*}=+26~{}\!{\rm mag}$ according to the survey depth of the Chinese Space Station Telescope (CSST; Gong et al., 2019). To select the candidate host galaxies, we adopt a $2\sigma$ confidence interval (CI) of the spatial localization error volumes, and convert the range of luminosity distance, $[D_{L;\min},D_{L;\max}]$, into a range of redshift, $[z_{\min},z_{\max}]$, using the relationship
$D_{L;\min}$ $=c(1+z_{\min})\int_{0}^{z_{\min}}\frac{1}{H^{-}(z^{\prime})}\mathrm{d}z^{\prime},$ (19a)
$D_{L;\max}$ $=c(1+z_{\max})\int_{0}^{z_{\max}}\frac{1}{H^{+}(z^{\prime})}\mathrm{d}z^{\prime},$ (19b)
where $H^{-}(z)$ and $H^{+}(z)$ are the minimum and maximum values of $H(z)$, respectively, given by the cosmological prior. For the prior, when we focus on the constraint on $H_{0}$, we choose uniform distributions of $h\in\mathcal{U}[0.6,0.8]$ and $\Omega_{M}\in\mathcal{U}[0.04,0.5]$ while fixing $w_{0}=-1$ and $w_{a}=0$; when we are interested in the constraint on the dark energy EoS, we choose uniform distributions of $w_{0}\in\mathcal{U}[-2,-0.5]$ and $w_{a}\in\mathcal{U}[-1,1]$ while fixing $h$ and $\Omega_{M}$ at their default values.
The distributions of the numbers of the candidate host galaxies are shown in Panel (c) of Figure 1. We can see that the typical number of candidate host galaxies for an EMRI detected by TianQin or TianQin I+II is about $10^{5}-10^{6}$ after accounting for the uncertainties in the cosmological parameters, and that the TianQin+LISA network can reduce the typical number to about $10^{3}-10^{4}$. In addition, to make the mock galaxy catalog more realistic, we assume that all galaxies with $z>0.6$ in the MultiDark galaxy catalog have photometric redshift (photo-$z$) errors which vary with the true redshifts $z_{\rm true}$ of the galaxies as
$\sigma_{{\rm photo-}z}(z_{\rm true})=\left\{\begin{array}[]{ll}0,&z_{\rm true}\leq 0.6,\\ 0.03(1+z_{\rm true}),&z_{\rm true}>0.6,\end{array}\right.$ (20)
(Dahlen et al., 2013; Zhou et al., 2021, 2022). The reason for setting the photo-$z$ errors to zero for galaxies with $z\leq 0.6$ is that spectroscopic surveys of ground-based telescopes can cover this redshift range (Cui et al., 2012; Dawson et al., 2013; Aghamousa et al., 2016).
To constrain the formation channel of EMRIs, we estimate the expected number of AGNs $\lambda_{i}$ within the spatial localization error volume $\Delta V_{i}$ of the $i$th EMRI according to the empirical bolometric AGN luminosity functions $\Phi(M^{*},z)$ (Hopkins et al., 2007), where $M^{*}$ represents the absolute magnitude of AGN, and then use this expectation to generate a Poisson distribution of random numbers. More quantitatively, we calculate the expectation $\lambda_{i}$ with
$\lambda_{i}=\Delta V_{i}\int_{-\infty}^{M_{{\rm limit},i}^{*}}\Phi(M^{*},z)\mathrm{d}M^{*},$ (21)
where
$M_{{\rm limit},i}^{*}\approx m_{\rm limit}^{*{\rm agn}}-25-5\lg\!\left(\!\frac{D_{L,i}}{1\!~{}{\rm Mpc}}\!\right)\!-BC$
represents the observational limiting absolute magnitude for the candidate host AGN of the $i$th EMRI at a distance of $D_{L,i}$, and $BC$ is a factor for bolometric correction (Duras et al., 2020). In this work, we set the limit magnitude of the AGN survey to $m_{\rm limit}^{*{\rm agn}}=+24.4~{}\!{\rm mag}$ according to the depth in the shallowest band of the CSST survey as well as taking into account the need for the spectroscopic identifications of AGNs (Gong et al., 2019). Taking the M1 model as an example, the distributions of the numbers of candidate host AGNs of the EMRIs for various detector configurations are shown in Panel (d) of Figure 1. For the same EMRIs, the numbers of candidate host AGNs are about $1\%$ of the numbers of the normal galaxies which are candidate hosts. This fraction is consistent with the AGN fraction among all galaxies.
4 Constraints on the cosmological parameters
In this section, we use the Bayesian framework described in Section 2.2 to calculate the posterior probability distributions of various cosmological parameters, including the standard $\Lambda$CDM model parameters $(H_{0},\Omega_{M})$ and the dark energy EoS parameters $(w_{0},w_{a})$. The essence of the computational process is fitting the $D_{L}-z$ relation with Markov Chain Monte-Carlo (MCMC) sampling. In the following, we adopt a total of 11 population models, from M1 to M12 except M11, to evaluate the capability of future EMRI observations in constraining the cosmological parameters.
4.1 Constraints on $H_{0}$ and $\Omega_{M}$
We first fit the $D_{L}-z$ relation, using the luminosity distances from EMRI observations and the redshifts extracted from galaxy surveys. The result is shown in Figure 2, and the corresponding constraints on $H_{0}$ and $\Omega_{M}$ are shown in Figure 3. Note that n this section we fix the parameters of the dark energy EoS to $w_{0}=-1$ and $w_{a}=0$.
Figure 2: Typical example of fitting the $D_{L}-z$ relation derived from the EMRI dark sirens which are detectable by the TianQin+LISA network. Here we use the fiducial EMRI population model. The horizontal colored serial squares represent the probability distribution of redshift (normalized by the maximum) derived from the observed candidate host galaxies, and the color hues from white to red show the probability of redshift. The luminosity distances shown here represent the mean values, and their errors ($1\sigma$ CI) are indicated by the black error bars. The solid green line represents the true cosmological model adopted by our simulations. The dotted-dashed red line represents the best-fit cosmological model, while the two shaded regions representing the $1\sigma$ and $2\sigma$ CIs.
Figure 3: Constraining errors of $h$ and $\Omega_{M}$ for TianQin (black), TianQin I+II (blue), and TianQin+LISA (red). The horizontal axis represents the different EMRI population models. The error bars correspond to a $1\sigma$ CI. The gray shade represents the $1\sigma$ equivalent error of the prior, and the green shade represents a error scale of $2\%$ ($10\%$) for $h$ ($\Omega_{M}$).
For the Hubble-Lemaître constant $H_{0}$, under the fiducial population model, i.e., the M1 model, the expected error is about $8.1\%$ for TianQin, and TianQin I+II can reduce this error to about $4.4\%$. Under the other population models of from M2 to M12, the error of $H_{0}$ varies considerably from model to model. The variations are mainly caused by the fact that the detection rates of EMRIs predicted by various models vary considerably (see Babak et al., 2017; Fan et al., 2020). In general, TianQin is able to achieve a precision in the range of about $3.4\%-9.6\%$, while that for TianQin I+II is expected to be in the range of about $1.5\%-7.2\%$.
For the fractional matter density parameter $\Omega_{M}$, in almost all population models the TianQin configuration cannot provide reasonable constraint (comparing the error scales corresponding to the priors and the posteriors), except for the M6, M7, and M12 models. But TianQin I+II can significantly improve the constraint on $\Omega_{M}$. More quantitatively, under the M6, M7, and M12 models, TianQin is likely to constrain $\Omega_{M}$ to a precision of about $46\%$, $33\%$, and $42\%$, respectively, whereas TianQin I+II would be able to constrain $\Omega_{M}$ to a precision of about $29\%$, $13\%$, and $11\%$, respectively.
If TianQin and LISA could form a network, i.e., TianQin+LISA, the constraints on $H_{0}$ and $\Omega_{M}$ are significantly improved in all population models, as is shown by the red error bars in Figure 3. For example, in the M1 model, the errors of $H_{0}$ and $\Omega_{M}$ can shrink to about $1.9\%$ and $20\%$, respectively. Under the other models, the precision of $H_{0}$ can be improved to about $0.4\%-7.1\%$, and the error of $\Omega_{M}$ would reduce to the range of about $13\%-43\%$. For clarity, we summarize the expected precisions on $H_{0}$ and $\Omega_{M}$ for various cases in Table I.
4.2 Constraints on $w_{0}$ and $w_{a}$
Figure 4: Same as Figure 3, but for $w_{0}$ and $w_{a}$. The green shades represent error scales of $\Delta w_{0}=0.1$ and $\Delta w_{a}=0.1$.
The errors on the dark energy EoS parameters $w_{0}$ and $w_{a}$ are shown in Figure 4. Here we fix $h=0.678$ and $\Omega_{M}=0.307$. We find that only $w_{0}$ can be effectively constrained and $w_{a}$ is not constrained in almost all population models except in the M7 and M12 models. The much better constraint on $w_{0}$ than on $w_{a}$ implies that inferring the deviation of the dark energy EoS with respect to the value $w=-1$ of the standard model requires much better EMRI GW detections than reconstructing the evolution of the dark energy EoS with redshift. Even under the M7 and M12 models, despite the fact that they have optimistic EMRI rates such that TianQin I+II and TianQin+LISA can detect a total number of more than one thousand EMRIs, still, a stringent constraint on $w_{a}$ cannot be obtained.
Table I: Expected relative errors on the $\Lambda$CDM cosmological model parameters $(H_{0},\Omega_{M})$ and the CPL dark energy model parameters $(w_{0},w_{a})$ for the 11 EMRI population models and assuming three different detector configurations: TianQin, TianQin I+II and TianQin+LISA. The bar “--” means no effective constraint.
EMRI Expected relative error
population TianQin TianQin I+II TianQin+LISA
model $\frac{\Delta H_{0}}{H_{0}}$ $\frac{\Delta\Omega_{M}}{\Omega_{M}}$ $|\frac{\Delta w_{0}}{w_{0}}|$ $\frac{\Delta w_{a}}{1}$ $\frac{\Delta H_{0}}{H_{0}}$ $\frac{\Delta\Omega_{M}}{\Omega_{M}}$ $|\frac{\Delta w_{0}}{w_{0}}|$ $\frac{\Delta w_{a}}{1}$ $\frac{\Delta H_{0}}{H_{0}}$ $\frac{\Delta\Omega_{M}}{\Omega_{M}}$ $|\frac{\Delta w_{0}}{w_{0}}|$ $\frac{\Delta w_{a}}{1}$
M1 8.1% -- 37% -- 4.4% 34% 12% -- 1.9% 20% 6.5% --
M2 7.9% -- 37% -- 2.5% 36% 12% -- 1.4% 19% 7.5% --
M3 5.6% -- 24% -- 5.3% 31% 9.6% -- 1.0% 12% 6.7% --
M4 6.1% -- 23% -- 4.2% -- 10% -- 2.9% 33% 7.2% --
M5 8.3% -- 39% -- 5.5% 47% 24% -- 2.1% 26% 6.6% --
M6 4.9% 46% 15% -- 2.9% 29% 7.3% -- 1.3% 15% 6.5% --
M7 2.4% 33% 8.7% -- 1.1% 13% 5.4% -- 0.4% 5.8% 4.1% 44%
M8 9.6% -- 48% -- 7.2% -- 43% -- 7.1% 46% 22% --
M9 5.4% -- 30% -- 3.1% 46% 14% -- 1.6% 23% 6.6% --
M10 5.9% -- 36% -- 4.0% -- 22% -- 1.8% 27% 7.8% --
M12 3.4% 42% 6.7% -- 1.5% 11% 5.4% 47% 0.4% 4.7% 4.2% 37%
More specifically, under the fiducial population model, the error of $w_{0}$ can reach about $37\%$ for TianQin, about $12\%$ for TianQin I+II, and about $6.5\%$ for TianQin+LISA. Under the other population models, i.e., from M2 to M12 except M11, the expected errors of $w_{0}$ can reach $6.7\%-48\%$ for TianQin, about $5.4\%-43\%$ for TianQin I+II, and about $4.2\%-22\%$ for TianQin+LISA. Similar to the constraints on $H_{0}$, the errors on $w_{0}$ also vary considerably across population models. Finally, we also summarize the expected constraint precisions on $w_{0}$ and $w_{a}$ for various cases in Table I.
5 Constraints on the formation channel of EMRIs
So far, we have assumed that each candidate host galaxy has the same probability of hosting an EMRI, whether it is a normal galaxy or an AGN. Now we drop the assumption and use the statistical framework described in Section 2.3 to constrain $f_{\rm agn}$, which is the fraction of EMRIs which reside in AGNs. To account for the random fluctuation of the AGN number density in the Universe, we use Monte-Carlo simulations to estimate the background distribution $P_{\rm bg}(\lambda)$ of the likelihood ratio $\lambda$ in Equation (18). We assume that $100$ EMRIs will be detected by the TianQin+LISA network (the number is expected to be reached with 5 years of detecting under most population models, except for the M5, M8, and M11 models, see Fan et al., 2020). For example, several background distributions for different $f_{\rm agn}$ are shown in Figure 5 as the dashed lines. In addition, to account for the randomness in simulated realizations, we also estimate the probability distribution $P(\lambda)$ through several times of independent simulations, as shown by the solid lines in Figure 5. We can see that the variation of $P(\lambda)$ is more significant than $P_{\rm bg}(\lambda)$ as $f_{\rm agn}$ increases, this is because the variation of $P(\lambda)$ is caused by a combination of changes in the value of $f_{\rm agn}$ and the numbers of candidate host AGNs $\{N_{{\rm agn},i}\}$ whereas the variation of $P_{\rm bg}(\lambda)$ is only caused by the change of $f_{\rm agn}$. Since the statistical significance mainly depends on the number of GW sources, we denote $N_{{\rm GW},3\sigma}^{\rm threshold}$ as the minimum number of detected EMRIs that is required to reject the null hypothesis at a significance level of $3\sigma$, and we study its dependence on the value of $f_{\rm agn}$.
Figure 5: Typical background distribution $P_{\rm bg}(\lambda)$ (dashed line) and probability distribution $P(\lambda)$ (solid line) of the likelihood ratio $\lambda$ in Equation (18) for 100 EMRIs detected by TianQin+LISA, under the M1 population model with fixed cosmology. The colors black, blue, cyan, and red represent four fractions of $f_{\rm agn}=0.1,0.3,0.6,1.0$, respectively.
Figure 6: Minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ of detected EMRI events required to testify the EMRI-AGN correlation with a $3\sigma$ significance. Here we have used the M1 population model. The black, blue, and red error bars and lines correspond to TianQin, TianQin I+II, and TianQin+LISA observations. The solid lines represent the case that both $H_{0}$ and $\Omega_{M}$ are fix to the true values, and the dashed lines with corresponding error bars represent the case in which both $H_{0}$ and $\Omega_{M}$ are freely variable within the prior. The error bars correspond to $1\sigma$ CIs.
Here, we take the M1 model as the fiducial population model again to estimate the required minimum number of detected EMRIs. When we convert the error of luminosity distance to the redshift error using Equations (19a) and (19b), we assume two cosmological conditions, namely: (i) “fixed cosmology” that the cosmological parameters are fixed to their true values, and (ii) “variable cosmology” that the cosmological parameters $H_{0}$ and $\Omega_{M}$ are freely variable within their priors. The latter condition is a conservative assumption, considering that there is a tension in current measurements of the Hubble constant (Freedman, 2017; Aghanim et al., 2020; Riess et al., 2021; Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022). The resulting dependence of the required minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ on the fraction $f_{\rm agn}$ is shown in Figure 6. We can see that $N_{{\rm GW},3\sigma}^{\rm threshold}$ decreases with increasing $f_{\rm agn}$ in all cases.
In the “fixed cosmology” condition, TianQin would need to detect at least $465$ EMRIs to be able to reject the null hypothesis and establish the EMRI-AGN correlation if $f_{\rm agn}=1$. Under the same condition, TianQin I+II needs to detect at least about $196$ EMRIs and TianQin+LISA needs to detect at least about $26$ EMRIs. In the “variable cosmology” condition, if $f_{\rm agn}=1$ the minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ required for TianQin, TianQin I+II, and TianQin+LISA are about $1650$, $584$, and $76$, respectively. The increase of $N_{{\rm GW},3\sigma}^{\rm threshold}$ in the case of variable cosmology is caused by an increase of the redshift range when transforming the luminosity distance range into the redshift space. For different cases, $N_{{\rm GW},3\sigma}^{\rm threshold}$ as a function of $f_{\rm agn}$ can be approximate as $N_{{\rm GW},3\sigma}^{\rm threshold}\propto f_{\rm agn}^{-1.2}$.
For the other EMRI population models, since the probability distributions of spatial localization errors of EMRIs detected by LISA and TianQin are not much different from the M1 model (Babak et al., 2017; Fan et al., 2020), the required minimum number of EMRI detections estimated in the M1 model also applies approximately to the other population models.
6 Discussions
6.1 Advantage of combining EMRI detections and AGN observations
Once either of the following two conditions is met, we can use AGN catalogs to extract the redshift information of the detected EMRIs (or at least assign higher weights to the AGNs): (i) the fraction of the EMRIs originate in AGNs, $f_{\rm agn}$, is determined to be close to $1$ by statistical analysis of the EMRI-AGN correlation; and (ii) a detected EMRI has a zero eccentricity at the plunge (see, e.g., Pan et al., 2021; Amaro-Seoane et al., 2023, for explanation). We note that LISA and TianQin can measure the eccentricities of EMRIs to a precision of $\Delta e\approx 10^{-6}$ (Babak et al., 2017; Fan et al., 2020), so they can differentiate the formation channels.
Assuming that all EMRIs come from AGNs, and using only the AGN catalog to look for the candidate hosts for EMRIs, we repeat the simulation shown in Section 3.2 and derive the posteriors for the cosmological parameters. More specifically, we consider the EMRIs detected by TianQin in the M1 population model as an example. The AGN catalog is based on the MultiDark catalog but we use the limiting magnitude of CSST to select observable AGNs. The result is shown in Figure 7. We find that the constraints on cosmological parameters are improved. More quantitatively, by switching from the galaxy catalog to the AGN catalog, the precision of $H_{0}$ improves from about $3.7\%$ to about $3.2\%$, and the precision of $w_{0}$ improves from about $25\%$ to about $13\%$. For a fair comparison, here we do not consider the redshift errors in either candidate host galaxies or AGNs.
Figure 7: Constraints on $(h,\Omega_{M})$ (left panel) and $(w_{0},w_{a})$ (right panel) by combining EMRI detections of TianQin with AGN catalogs. The eccentricities at plunges are all set to zero, and the other parameters are derived from the M1 model. The blue contours represent the constraints derived from candidate host galaxies and the red contours represent the constraints from candidate host AGNs, respectively. The two levels of contour correspond to $1\sigma$ and $2\sigma$ CIs. The cyan crosses represent the true values of the cosmological parameters used in our simulations.
These improvements mainly benefit from the fact that AGNs constitute only about $(1-10)\%$ of all galaxies (e.g., compare the galaxy and AGN luminosity functions in Dahlen et al., 2005; Hopkins et al., 2007). When $f_{\rm agn}$ is close to $1$ or the detected EMRIs have zero eccentricities at plunge, we can derive more precisely the redshifts of EMRIs with the help of AGN catalogs. These results also imply that circular EMRIs are better than those eccentric ones in constraining cosmological parameters.
In addition, we would like to point out that EMRI GW sources may be accompanied by observable EM signals (Wang et al., 2019b, 2022b; Li et al., 2023; Amaro-Seoane et al., 2023). As long as the EM signals from EMRI sources do exist, current and planning time-domain survey telescopes, such as the Wide Field Survey Telescope (WFST) (Wang et al., 2023; Huang et al., 2024), the Multi-channel Photometric Survey Telescope (Mephisto) (Yuan et al., 2020), Vera Rubin Observatory (also known as LSST; Ivezić et al., 2019), the Einstein Probe (Yuan et al., 2015), Athena (Nandra et al., 2013; McGee et al., 2020), the Five-Hundred Aperture Spherical Radio Telescope (FAST) (Nan et al., 2011), the Square Kilometre Array (SKA) (Janssen et al., 2015), and so on, could provide promising searches for the EM signals. Once the EM signals of EMRIs are observed, they could help us narrow down the pools of the candidate host galaies or even uniquely identify the host galaxies of the EMRIs. This will allow us to not only improve the precisions of the cosmological parameters by more accurately measuring the redshifts of the EMRIs but also enhance our understanding for the formation channel of the EMRIs.
6.2 Effects of eccentricity on the spatial localization of EMRIs
Since we have set the eccentricities of the EMRIs to be uniformly distributed in the range of 00 to $0.2$ when simulating the EMRI detections, however, the EMRIs originated in AGNs are expected to have a zero eccentricity at plunge (Pan et al., 2021; Amaro-Seoane et al., 2023). So that if there is a fraction of $f_{\rm agn}$ of EMRIs that do originate in AGNs, the values of eccentricities of EMRIs may affect the measurement precisions of their source parameters. Here, we analyze the effect of eccentricities of EMRIs at plunges on their spatial localization errors, the result is shown in Figure 8.
From Figure 8, we can see that the effect of EMRIs’ eccentricities at plunges on their spatial localization errors (sky position and distance) are not significant, and the error increases with the decrease of the eccentricity very slightly. This implies that both the expected errors on the cosmological parameters presented in Section 4 and the required minimum number of EMRI detections for establishing the EMRI-AGN correlation presented in Section 5 will not be significantly changed.
Figure 8: Variation of the spatial localization errors of TianQin EMRIs caused by their eccentricities at plunges. The gray (red) violinplots represent the probability distributions for the sky localizations ($D_{L}$ estimations) of EMRIs, and the gray (red) horizontal dashed line represents the median of the first violinplot.
6.3 Prospect of clarifying the Hubble tension
The cosmological standard model—flat-$\Lambda$CDM, is challenged by the Hubble tension (Freedman, 2017; Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022). The measurement of $H_{0}$ obtained by the flat-$\Lambda$CDM model plus cosmic microwave background observations is $67.4\pm 0.5~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$, with a relative error of about $0.7\%$ (Aghanim et al., 2020), and the $H_{0}$ measurement obtained by the cosmic distance ladder calibrations and type Ia supernova observations is $73.2\pm 1.3~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$, with a relative error of about $1.8\%$ (Riess et al., 2021). So if we want to clarify the Hubble tension by a new independent measurement, the independently measured $H_{0}$ needs to be better than a precision of about $2\%$.
We can see from Table I that TianQin does not achieve a precision of 2% in constraining $H_{0}$ under all EMRI population models. As for TianQin I+II, it achieves a precision better than $2\%$ only under the population models with optimistic EMRI rates, such as the M7 and M12 models. However, the multi-detector network formed by LISA and TianQin, i.e., TianQin+LISA, can much better constrain $H_{0}$. For example, under most of the population models, except for the M8 and M11 models, the TianQin+LISA network can constrains $H_{0}$ to a precision of near or better than $2\%$. Therefore, the results of our work support a joint observation of GWs by TianQin and LISA.
7 Conclusion
In this work, we have investigated the potential of using EMRIs as dark sirens to constrain the cosmological parameters as well as the formation channel of EMRIs. We adopted 11 population models presented in Babak et al. (2017) to simulate the detected EMRI GW event catalogs.
(i) To constrain the cosmological parameters, we derive the redshift probability distributions of detected EMRIs by matching their spatial localization error volumes with galaxy survey catalogs, and then evaluate the posterior probability distributions of the cosmological parameters via a Bayesian analytical framework which combines the luminosity distance and redshift information. Our results show that precise constraints on the Hubble-Lemaître constant $H_{0}$ and the dark energy EoS parameter $w_{0}$ can be obtained from EMRIs. Under the fiducial population model M1, the precision of $H_{0}$ can reach about $8.1\%$ for TianQin, about $4.4\%$ for TianQin I+II, and about $1.9\%$ for the TianQin+LISA network; and the precision of $w_{0}$ can reach about $37\%$ for TianQin, about $12\%$ for TianQIn I+II, and about $6.5\%$ for the TianQin+LISA network. Additionally, under the most optimistic EMRI population model, TianQin can constrain $H_{0}$ and $w_{0}$ to a precision better than about $3\%$ and about $10\%$, and the TianQin+LISA network can constrain $H_{0}$ and $w_{0}$ to a precision better than about $1\%$ and about $5\%$, respectively.
(ii) To constrain the formation channel of EMRIs, we adopt an AGN luminosity function to randomly generate the numbers of candidate host AGNs of the detected EMRIs, and then use Poisson distributions to statistically test the spatial correlation between the EMRIs and AGNs. Under the fiducial population model, M1, we mainly find that the minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ of the EMRIs required to establish the EMRI-AGN correlation decreases with increasing $f_{\rm agn}$. If $f_{\rm agn}=1$, for the three detector configurations of TianQin, TianQin I+II, and TianQin+LISA, the values of $N_{{\rm GW},3\sigma}^{\rm threshold}$ respectively are about $465$, $196$, and $26$ when the cosmological parameters are fixed. Under the condition that the cosmological parameters vary freely within the priors, the values of $N_{{\rm GW},3\sigma}^{\rm threshold}$ become significantly larger, all by a factor of about three or a bit larger. For other population models, $N_{{\rm GW},3\sigma}^{\rm threshold}$ is similar. Moreover, after we confirm that a large fraction of EMRIs do indeed originate in AGNs, the combination of the EMRIs and AGN catalogs can significantly improve the constraints on the cosmological parameters.
The cosmological constraints reported in this work appear to be somewhat weaker than those of LISA alone reported in Laghi et al. (2021) (also see MacLeod & Hogan, 2008). For example, under the M1, M5, and M6 models, LISA with a four year mission is expected to constrain $H_{0}$ and $w_{0}$ to higher precisions than TianQin or TianQin I+II. This is mainly because the most sensitive band of TianQin is slightly higher than that of LISA, which makes the EMRI detection rate of TianQin much lower than LISA (Babak et al., 2017; Fan et al., 2020). However, we would like to point out that currently there is large uncertainty in the mass function of MBHs especially at the low mass end (Greene & Ho, 2007; Barausse, 2012; Gallo & Sesana, 2019; He et al., 2024). The GW signals radiated by the EMRIs that are formed by lighter MBHs, will have higher frequencies and are more easily detected by TianQin, thus enhancing the total EMRI detection rate of TianQin.
Our analytical results of constraining the EMRI-AGN correlation demonstrate for the first time the prospects of using space-borne GW detectors to statistically infer the formation channel of EMRIs. Information about some intrinsic parameters of EMRI, such as the eccentricity at plunge (Pan et al., 2021) and the spin of MBH (Burke et al., 2020), can help us trace the formation channel of the EMRI, but an unbiased extraction of the intrinsic parameters relies heavily on an accurate EMRI GW waveform model (Chua et al., 2017; Amaro-Seoane, 2018; Torres-Orjuela et al., 2021). The effectiveness of the statistical framework proposed in this work mainly relies on the spatial localization accuracies of EMRIs and the completenesses of AGN catalogs. Since the localization accuracies are mainly determined by the SNRs of the EMRIs and the completenesses of the AGN catalogs are mainly determined by the depths of galaxy surveys, the accuracy of the waveform model is not as essential in our framework as in other studies. Therefore, our statistical framework can serve as an important supplementary means of inferring the formation channel of EMRIs.
Acknowledgements
The authors would like to thank Linhua Jiang, Yuming Fu, Zhenwei Lyu and Tieguang Zi for very helpful discussions, and thank the Theoretical Astrophysical Observatory (TAO; https://tao.asvo.org.au/tao/) for providing access to the download of the mock galaxy catalogs. L.-G. Zhu is funded by China Postdoctoral Science Foundation (Grant No. 2023M740113), X. Chen is supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203002) and the National Natural Science Foundation of China (Grant No. 11991053), Y.-M. Hu is supported by the the Natural Science Foundation of China (Grants No. 12173104 and No. 12261131504), and Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001), and J.-d. Zhang is supported by the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515030116). This work also supported by Key Laboratory of TianQin Project (Sun Yat-sen University), Ministry of Education. The computations in this work was performed on the High Performance Computing Platform of the Centre for Life Science, Peking University.